18. A body starting from rest acquires a velocity of 18 cm/s in 3 s. If the body has a mass of 15 g. What force is exerted on the body?


1
Expert's answer
2021-09-28T11:14:38-0400

The acceleration and force (by Newton's second law):


"a=\frac vt,\\\space\\
F=ma=m\frac vt=0.015·\frac{0.18}{3}=9·10^{-4}\text{ N}."